OpenStudy (anonymous):

A ball is thrown vertically upward. After t seconds, its height h (in feet) is given by the function h(t) =96t-16t^2 What is the maximum height that the ball will reach? h (in feet) is given by the function

OpenStudy (alekos):

maximum height occurs when h'(t) = 0 and then solve for t

OpenStudy (anonymous):

@Er.Mohd.AMIR help plz

OpenStudy (er.mohd.amir):

d/dt(h)=96-16*2t and 96-32t=0 32t=96 t=96/32 t=3 sec

OpenStudy (alekos):

Where's Missy? Just gets the answer and disappears with no thanks at all!!

OpenStudy (anonymous):

@alekos I always thank for the help I receive. I was booted off yesterday and just now was able to resign in. that is why I just closed the question.. once again THANKS for the help!!!

OpenStudy (alekos):

GOOD!!!